OpenStudy (anonymous):

The formula for converting degrees Fahrenheit (F) to degrees Celsius (C) is C = 5/9 (F – 32). Find C for F = 5°. Okay, I have tried this problem over 3 times and get a different answer each time, I must be doing something wrong. Can you guys help me?

OpenStudy (anonymous):

\[\frac{5}{9}(5-32)=-15\]

OpenStudy (anonymous):

5/9(-27)= 5*(-27) / 9 =-15

OpenStudy (anonymous):

Oh ok, that explains it, I was doing the problem totally wrong. Thank you guys!

OpenStudy (anonymous):

Remember, if you're using a calculator, you have to do it so the order of operations is not disturbed.